OpenStudy (anonymous):

solve the seperable equation: dx/dt=3xt^2

OpenStudy (anonymous):

this is a first-order linear ordinary differential equation, and the differential equation solution is x(t) = c_1 e^(t^3)

OpenStudy (jamesj):

Separating variables: dx / x = 3t^2 dt Now integrate both sides and you have ln x = t^3 + C Thus ....

OpenStudy (anonymous):

thank you very much!

OpenStudy (anonymous):

jamesj, got a second?

OpenStudy (jamesj):

sure ... chat?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ln(x)=t^3+C or x=Cexp(t^3)